OpenStudy (anonymous):

Can someone please answer this!? I will give out medals!! A student converted 480 ft to inches as follows. What did the student do wrong? What is the correct answer? 1 ft/12 in. = x/480 ft

OpenStudy (abmon98):

1/12=x/480, 480=12x , x=480/12, x=40 and this is incorrect 480 feet*12=5760 inches

OpenStudy (anonymous):

So what should i write? @Abmon98

OpenStudy (abmon98):

you could write my working as proove that the student is wrong